# Retrieval-Augmented Generation in Biomedicine: A Survey of Technologies, Datasets, and Clinical Applications

### v1

@@ Introduction
Clinical research application tools are designed to support clinical studies by enhancing efficiency, leveraging knowledge retrieval and synthesis to accelerate the development of medical knowledge and therapeutic innovations. 68.
Recent literature highlights several innovative RAG applications in clinical research leveraging different retrieval approaches. PodGPT 94 implements the BAAI embedding model for scientific literature retrieval informing research design and methodology development. LITURAt 169 utilizes the Entrez API for PubMed retrieval, supporting scientific data analysis in clinical research through relevant literature identification. These applications showcase RAG’s ability to optimize and improve the clinical research process through intelligent knowledge retrieval, highlighting its potential as a valuable tool for both research and education, and ultimately accelerating the translation of research findings into clinical practice.
@@
Recent research indicates that RAG systems can be both maliciously and unintentionally manipulated to produce unreliable decisions and potentially harm humans 53, 45. Such vulnerabilities are particularly unacceptable in the biomedical domain due to the potential for severe consequences.
Several dimensions of trustworthiness in biomedical RAG systems raise critical concerns. Robustness refers to the system’s ability to exclude malicious content introduced by attackers or filter inadvertent misinformation contained in web corpora that might contaminate the knowledge source 66. Although certain re-ranking methodologies described in Section Reranker can partially mitigate these issues, significant challenges remain in biomedical RAG applications. Explainability requires a comprehensive understanding of user intent to generate reliable responses aligned with ground-truth documents. A notable challenge emerges from diverse linguistic styles in input information, creating misalignment between user queries and knowledge sources. For instance, in medical QA systems, users may formulate queries in conversational language while knowledge sources are typically based on academic publications 122. This linguistic discrepancy poses substantial challenges for biomedical information retrieval, consequently diminishing the performance of biomedical RAG systems. Fairness dictates that trustworthy RAG systems must avoid discriminatory outputs during the generation process. Due to the uneven distribution of demographic attributes in training corpora 81, which may be present in both the pre-training data of LLMs and the knowledge sources utilized in RAG systems, biomedical RAG applications are susceptible to intrinsic biases. Addressing the fairness concerns also represents an important direction for future research.
@@
Among our surveyed publications, English remains the dominant language in all Biomedical RAG systems. This predominance is primarily attributed to the dominance of English in public biomedical knowledge repositories and medical task datasets. The performance of biomedical RAG systems in other languages, particularly those with limited resources, remains inadequately explored 161, 101. This fact presents two significant challenges in the development of multi-lingual biomedical RAG systems.
First, training LLMs specialized in low-resource languages is inherently difficult, which diminishes the generative capabilities of Biomedical RAG in other linguistic contexts. Second, the scarcity of available knowledge sources in non-dominant languages significantly constrains the augmentation potential afforded by knowledge retrieval mechanisms. Although some researchers have attempted to address this limitation through synthetic data generated by LLMs or translated from resource-rich languages 157, consistent and sustained efforts are still required to overcome this challenge in multilingual biomedical information retrieval and generation.
@@
Another challenge lies in extending retrieval capabilities beyond textual data to encompass multi-modalities highly prevalent in the biomedical domain, including images (e.g. X-ray 104, 170, 134), time series (e.g. ECG 239), sound recordings (e.g. lung sounds recorded 57), and video 77. Among the studies we identified, the overwhelming majority focus exclusively on text-based retrieval, with only a limited subset addressing image-based retrieval, while other modalities remain largely unexplored. Multi-modal data presents several intractable challenges for retrieval systems. For instance, in medical imaging, although we can utilize a visual encoder model as a retriever, high intra-class variance significantly complicates image retrieval processes 249. Specifically, identical medical conditions can manifest with remarkable heterogeneity across visual representations; for instance, images of histologically identical tumors may exhibit substantially different morphological characteristics when visualized across different patients or at varying stages of disease progression. Attention also should be paid to developing cross-modal RAG in the Biomedical domain and improving its accuracy and trustworthiness in the future.
@@
The application of RAG in the biomedical domain faces significant limitations due to computing resource constraints, which restrict its utility in practical scenarios. Within the RAG framework, LLMs play a crucial role in contextual understanding and human-like response generation based on retrieved knowledge 53. However, these models typically require high-performance computing clusters to operate effectively. For instance, the minimum deployment unit of DeepSeek V3’s decoding stage consists of 40 nodes with 320 GPUs 43.
Empirical research consistently demonstrates that LLMs with larger parameter counts generally exhibit superior capacity for understanding and generation tasks 27, 85. Nevertheless, in real-world healthcare environments such as hospitals, maintaining high-performance computing infrastructure is often impractical and cost-prohibitive. Several promising approaches have emerged to address these constraints. One solution involves developing smaller, distilled LLMs derived from their larger counterparts 235. Models such as QWQ 175 and Mistral Small 3.1 156 demonstrate competitive performance despite having significantly reduced parameter counts. Applying Portable LLMs 112 as generation models is also a promising method. Another viable approach incorporates RAG systems that combine local knowledge bases with external LLM API services. This hybrid configuration presents an acceptable compromise when computing resources are limited, although it introduces potential privacy concerns, as local data must be transmitted to commercial providers for response generation.
@@
Privacy concerns in biomedical RAG systems exist throughout both retrieval and generation processes. During knowledge retrieval, reference documents frequently contain patients’ clinical histories, which may include personally identifiable information that individuals would prefer to keep confidential 153. Such information handling might not comply with stringent regulatory frameworks such as the General Data Protection Regulation 38 and the Health Insurance Portability and Accountability Act 2. Furthermore, the generation phase presents additional challenges, particularly when utilizing commercial LLMs for biomedical RAG development. This approach necessitates the transmission of sensitive data across the internet, significantly increasing the risk of data breaches and unauthorized access. Consequently, developing biomedical RAG systems with robust privacy-preserving mechanisms remains a critical direction for future research endeavors.

### v2

@@ Introduction
Clinical research application tools are designed to support clinical studies by enhancing efficiency, leveraging knowledge retrieval and synthesis to accelerate the development of medical knowledge and therapeutic innovations. 230.
Recent literature highlights several innovative RAG applications in clinical research leveraging different retrieval approaches. PodGPT 34 implements the BAAI embedding model for scientific literature retrieval informing research design and methodology development. LITURAt 62 utilizes the Entrez API for PubMed retrieval, supporting scientific data analysis in clinical research through relevant literature identification. These applications showcase RAG’s ability to optimize and improve the clinical research process through intelligent knowledge retrieval, highlighting its potential as a valuable tool for both research and education, and ultimately accelerating the translation of research findings into clinical practice.
@@
Recent research indicates that RAG systems can be both maliciously and unintentionally manipulated to produce unreliable decisions and potentially harm humans 1, 231. Such vulnerabilities are particularly unacceptable in the biomedical domain due to the potential for severe consequences.
Several dimensions of trustworthiness in biomedical RAG systems raise critical concerns. Robustness refers to the system’s ability to exclude malicious content introduced by attackers or filter inadvertent misinformation contained in web corpora that might contaminate the knowledge source 181. Although certain re-ranking methodologies described in Section Reranker can partially mitigate these issues, significant challenges remain in biomedical RAG applications. Explainability requires a comprehensive understanding of user intent to generate reliable responses aligned with ground-truth documents. A notable challenge emerges from diverse linguistic styles in input information, creating misalignment between user queries and knowledge sources. For instance, in medical QA systems, users may formulate queries in conversational language while knowledge sources are typically based on academic publications 139. This linguistic discrepancy poses substantial challenges for biomedical information retrieval, consequently diminishing the performance of biomedical RAG systems. Fairness dictates that trustworthy RAG systems must avoid discriminatory outputs during the generation process. Due to the uneven distribution of demographic attributes in training corpora 2, which may be present in both the pre-training data of LLMs and the knowledge sources utilized in RAG systems, biomedical RAG applications are susceptible to intrinsic biases. Addressing the fairness concerns also represents an important direction for future research.
@@
Among our surveyed publications, English remains the dominant language in all Biomedical RAG systems. This predominance is primarily attributed to the dominance of English in public biomedical knowledge repositories and medical task datasets. The performance of biomedical RAG systems in other languages, particularly those with limited resources, remains inadequately explored 232, 233. This fact presents two significant challenges in the development of multi-lingual biomedical RAG systems.
First, training LLMs specialized in low-resource languages is inherently difficult, which diminishes the generative capabilities of Biomedical RAG in other linguistic contexts. Second, the scarcity of available knowledge sources in non-dominant languages significantly constrains the augmentation potential afforded by knowledge retrieval mechanisms. Although some researchers have attempted to address this limitation through synthetic data generated by LLMs or translated from resource-rich languages 234, consistent and sustained efforts are still required to overcome this challenge in multilingual biomedical information retrieval and generation.
@@
Another challenge lies in extending retrieval capabilities beyond textual data to encompass multi-modalities highly prevalent in the biomedical domain, including images (e.g. X-ray 203, 235, 236), time series (e.g. ECG 152), sound recordings (e.g. lung sounds recorded 237), and video 238. Among the studies we identified, the overwhelming majority focus exclusively on text-based retrieval, with only a limited subset addressing image-based retrieval, while other modalities remain largely unexplored. Multi-modal data presents several intractable challenges for retrieval systems. For instance, in medical imaging, although we can utilize a visual encoder model as a retriever, high intra-class variance significantly complicates image retrieval processes 239. Specifically, identical medical conditions can manifest with remarkable heterogeneity across visual representations; for instance, images of histologically identical tumors may exhibit substantially different morphological characteristics when visualized across different patients or at varying stages of disease progression. Attention also should be paid to developing cross-modal RAG in the Biomedical domain and improving its accuracy and trustworthiness in the future.
@@
The application of RAG in the biomedical domain faces significant limitations due to computing resource constraints, which restrict its utility in practical scenarios. Within the RAG framework, LLMs play a crucial role in contextual understanding and human-like response generation based on retrieved knowledge 1. However, these models typically require high-performance computing clusters to operate effectively. For instance, the minimum deployment unit of DeepSeek V3’s decoding stage consists of 40 nodes with 320 GPUs 240.
Empirical research consistently demonstrates that LLMs with larger parameter counts generally exhibit superior capacity for understanding and generation tasks 241, 242. Nevertheless, in real-world healthcare environments such as hospitals, maintaining high-performance computing infrastructure is often impractical and cost-prohibitive. Several promising approaches have emerged to address these constraints. One solution involves developing smaller, distilled LLMs derived from their larger counterparts 243. Models such as QWQ 160 and Mistral Small 3.1 40 demonstrate competitive performance despite having significantly reduced parameter counts. Applying Portable LLMs 244 as generation models is also a promising method. Another viable approach incorporates RAG systems that combine local knowledge bases with external LLM API services. This hybrid configuration presents an acceptable compromise when computing resources are limited, although it introduces potential privacy concerns, as local data must be transmitted to commercial providers for response generation.
@@
Privacy concerns in biomedical RAG systems exist throughout both retrieval and generation processes. During knowledge retrieval, reference documents frequently contain patients’ clinical histories, which may include personally identifiable information that individuals would prefer to keep confidential 245. Such information handling might not comply with stringent regulatory frameworks such as the General Data Protection Regulation 246 and the Health Insurance Portability and Accountability Act 247. Furthermore, the generation phase presents additional challenges, particularly when utilizing commercial LLMs for biomedical RAG development. This approach necessitates the transmission of sensitive data across the internet, significantly increasing the risk of data breaches and unauthorized access. Consequently, developing biomedical RAG systems with robust privacy-preserving mechanisms remains a critical direction for future research endeavors.